OpenStudy (anonymous):

.show that the following are spanning sets for R3 {(1,1,1)transpose (1,1,0)transpose (1,0,0)transpose}

OpenStudy (anonymous):

any coordinate in R3 (a,b,c) can be given in terms of linear combination of the three vectors...

OpenStudy (anonymous):

but how to find linear combination

OpenStudy (anonymous):

x + y + z = a x + y = b x = c y = b - c z = a - b hence linear combination is (a - b)(1,1,1) + (b-c) (1,1,0) + c(1,0,0) for any (a,b,c)